A triangle is made from a rope. The sides of the triangle are 25 cm, 11 cm and 31 cm. What will be the approx area of the square made from the same rope?
A) 67 cm2
B) 280 cm2
C) 134 cm2
D) None of the above

please give solution for dis problem

Answer by ewatrrr(24785)


Hi
A triangle is made from a rope. The sides of the triangle are 25 cm, 11 cm and 31 cm
rope is 67 cm long: 67/4 = 16.75cm for each side of square
A = s^2 = 280.05625 cm^2